a) Computer 2 to the power of 2007, then take (mod 41) of the whole thing

b) Computer 5 to the power of 50,000, then take (mod 31) of the whole thing
Answer #1

Few things to remember :

ax+y(mod n) =ax(mod n). ay(mod n)

axy (mod n) = (ax(mod n)) y

Now if ax(mod n) = b then,

(ax(mod n)) y = (b (mod n)) y = (by (mod n))

If you have any questions comment down.. Also please upvote thanks..